OpenStudy (anonymous):

how do i simplify 1/sqrt(x+h) + sqrt(x) (limits and difference quotient)

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x+h} +\sqrt{x}}\]

OpenStudy (anonymous):

book answer says its simplified to\[\frac{ 1 }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

rationalize

OpenStudy (anonymous):

main problem: find slope of function f(x) = sqrt(x)

OpenStudy (anonymous):

what do i rationalize? x from the bottom?